OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1=18 and a15=-38? a.-142 b.-138 c.-134 d.-130

OpenStudy (anonymous):

\[a _{1}=18 \] and \[a _{15}=-38 \]

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@Sloth7 how did you get it?

OpenStudy (anonymous):

18 + 14d = -38 14d = -56 d = -4 18 + 40*-4 = 18 - 160 = -142

OpenStudy (anonymous):

Thank You for your help, it is greatly appreciated @Sloth7

OpenStudy (luigi0210):

Nice work @Sloth7

OpenStudy (anonymous):

TY